Miller Toy Company manufactures a plastic swimming pool at its Westwood Plant. The plant has been experiencing problems as shown by its June contribution format income statement below:

Budgeted Actual
  Sales (4,000 pools) $ 275,000    $ 275,000   
    
  Variable expenses:      
     Variable cost of goods sold* 74,720    90,040   
     Variable selling expenses 27,000    27,000   
    
  Total variable expenses 101,720    117,040   
    
  Contribution margin 173,280    157,960   
    
  Fixed expenses:      
     Manufacturing overhead 68,000    68,000   
     Selling and administrative 93,000    93,000   
    
  Total fixed expenses 161,000    161,000   
    
  Net operating income (loss)    $ 12,280    $ (3,040)
    
*Contains direct materials, direct labor, and variable manufacturing overhead.

Janet Dunn, who has just been appointed general manager of the Westwood Plant, has been given instructions to “get things under control.” Upon reviewing the plant’s income statement, Ms. Dunn has concluded that the major problem lies in the variable cost of goods sold. She has been provided with the following standard cost per swimming pool:

Standard Quantity or Hours Standard Price
or Rate
Standard Cost
  Direct materials    4.3 pounds $ 2.90 per pound $ 12.47   
  Direct labor    0.6 hours $ 8.40 per hour    5.04   
  Variable manufacturing overhead    0.3 hours* $ 3.90 per hour    1.17   
    
  Total standard cost $ 18.68   
    
*Based on machine-hours.
     During June the plant produced 4,000 pools and incurred the following costs:
a.

Purchased 22,200 pounds of materials at a cost of $3.35 per pound.

b.

Used 17,000 pounds of materials in production. (Finished goods and work in process inventories are insignificant and can be ignored.)

c. Worked 3,000 direct labor-hours at a cost of $8.10 per hour.
d.

Incurred variable manufacturing overhead cost totaling $6,450 for the month. A total of 1,500 machine-hours was recorded.

It is the company’s policy to close all variances to cost of goods sold on a monthly basis.
Required:
1. Compute the following variances for June:
a.

Materials price and quantity variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

         

b.

Labor rate and efficiency variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

           

c.

Variable overhead rate and efficiency variances. (Do not round your intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

           

2.

Summarize the variances that you computed in (1) above by showing the net overall favorable or unfavorable variance for the month. (Input all values as positive amounts. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

       

3.

Pick out the two most significant variances that you computed in (1) above. (You may select more than one answer. Single click the box with a check mark for correct answers and double click to empty the box for the wrong answers.)

Materials price variance
Labor efficiency variance
Variable overhead efficiency variance
Labor rate variance
Variable overhead rate variance
Materials quantity variance
Answer #1

Material Price Variance = (Actual Price - Standard Price) x Actual Quantity Used
= ($3.35 - $2.90) x 17000
= $7650 (U)

It can be calculated for Actual Quantity purchased as well
Material Price Variance = (Actual Price - Standard Price) x Actual Quantity Used
= ($3.35 - $2.90) x 22200
= $9990 (U)

Material Quantity Variance = (Actual Quantity - Standard Quantity) x Standard Price
= (17000 - 17200) x $2.90 = $580 (F)
Standard Quantity = 4000 x 4.3 = 17200

Labor Rate variance = (Actual Rate - Standard rate) x Actual hours
= ($8.10 - $8.40) x 3000 = $900 (F)

Labor efficiency variance = (Actual Hours - standard hours) x Standard rate
= (3000 - 2400) x $8.40 = $5040 (U)
Standard hours = 4000 x 0.6 = 2400

Variable overhead Rate variance = Actual Overhead - Standard rate x Actual hours
= $6450 - $3.90 x 1500 = $600 (U)

Variable Overhead efficiency variance = (Actual Hours - standard hours) x Standard rate
= (1500 - 1200) x $3.90 = $1170 (U)
Standard hours = 4000 x 0.3 = 1200

2.

Material Price Variance $          7,650 U
Material Quantity Variance $            -580 F
Labor Rate Variance $            -900 F
Labor Efficiency Variance $          5,040 U
Variable Overhead Rate Variance $              600 U
Variable Overhead Efficiency Variance $          1,170 U
$        12,980 U

3.
Most significant variances are Material Price Variance and Labor efficiency Variance